OpenStudy (anonymous):

Which of the following is an example of the additive identity? a + (-a) = 0 a + 0 = 0 a + 0 = a

OpenStudy (anonymous):

the additive identity is an element that can be added to any other without changing its value -- so in this case, \(a+0=a\)

OpenStudy (marihelenh):

The multiplicative identity is a*1=a. This means that when it is multiplied by one, it is still itself. So, for the additive identity, you need to get a is equal to itself, so a+0=a.

OpenStudy (anonymous):

\(a+(-a)=0\) is a demonstration of the additive inverse property, namely that there is an element \(-a\) for \(a\) that reduces it to the additive identity when added

OpenStudy (anonymous):

@oldrin.bataku @marihelenh thanks